A company assigns each of its employees an ID code that consists of three uppercase letters followed by two digits from 0 through 9.

a.   How many employee codes does the company have available?

b.   How many employee codes begin with HOW, WHO or WHY?
Answer #1

a) Number of employee codes available = 26*26*26*10*10 = 1757600

b) First three letter are fixed. So,

Number of employee codes begin with HOW, WHO or WHY = 10*10 + 10*10 + 10*10 = 300